OpenStudy (anonymous):

What is the 12th term of the Fibonacci Sequence?

OpenStudy (anonymous):

I started from after the 9th because that is when my chart ends. I just want to confirm my answer.

OpenStudy (amistre64):

a F sequence gets the next term by adding the 2 previous terms together

OpenStudy (anonymous):

Yes. I got 144 by using the steps 21 + 34 = 55 34 + 55 = 89 55 + 89 = 144. Is this correct?

OpenStudy (amistre64):

might be easier to keep track of if you create a column 0 1 1 2 3 5 8 13 21 34 55 89 144 233 377 ......

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

ok. Thank you.